# Pay Attention to Real World Perturbations! Natural Robustness Evaluation in Machine Reading Comprehension

### v1

@@ Abstract
As neural language models achieve human-comparable performance on Machine Reading Comprehension (MRC) and see widespread adoption, ensuring their robustness in real-world scenarios has become increasingly important. Current robustness evaluation research, though, primarily develops synthetic perturbation methods, leaving unclear how well they reflect real life scenarios. Considering this, we present a framework to automatically examine MRC models on naturally occurring textual perturbations, by replacing paragraph in MRC benchmarks with their counterparts based on available Wikipedia edit history. Such perturbation type is natural as its design does not stem from an arteficial generative process, inherently distinct from the previously investigated synthetic approaches. In a large-scale study encompassing SQuAD datasets and various model architectures we observe that natural perturbations result in performance degradation in pre-trained encoder language models. More worryingly, these state-of-the-art Flan-T5 and Large Language Models (LLMs) inherit these errors. Further experiments demonstrate that our findings generalise to natural perturbations found in other more challenging MRC benchmarks. In an effort to mitigate these errors, we show that it is possible to improve the robustness to natural perturbations by training on naturally or synthetically perturbed examples, though a noticeable gap still remains compared to performance on unperturbed data.
@@ 1 Introduction
Transformer-based pre-trained language models demonstrate remarkable efficacy in addressing questions based on a given passage of text, a task commonly referred to as Machine Reading Comprehension (MRC) (Devlin et al., 2019; Brown et al., 2020; He et al., 2021; Wei et al., 2022; Touvron et al., 2023; OpenAI et al., 2024b). Despite these advancements, high-performing MRC systems are also known to succeed by relying on shortcuts in benchmark datasets rather than truly demonstrating understanding of the passage, thereby lacking robustness to various types of test-time perturbations (Ho et al., 2023; Schlegel et al., 2023; Levy et al., 2023).
Evaluating models’ resilience to textual perturbations during inference aids in identifying adversarial instances that highlight their shortcut behavior and provides insights into mitigating these shortcuts (Ho et al., 2023). While numerous synthetic perturbation approaches have been explored and reveal the vulnerabilities of MRC models to various linguistic challenges (Ribeiro et al., 2018; Jiang and Bansal, 2019; Welbl et al., 2020; Tan et al., 2020; Tan and Joty, 2021; Schlegel et al., 2021; Cao et al., 2022; Tran et al., 2023), a serious concern is that these carefully designed perturbations might not necessarily appear in real-world settings. Consequently, this poses a risk of neglecting the weaknesses of reading comprehension systems to real challenges when deployed in practical scenarios, thus potentially hindering the improvement of their reliability in practical applications.
@@
To counteract this issue, in this paper, we develop a framework to inject textual changes that arise in real-world conditions into MRC datasets and audit how well contemporary language models perform under such perturbations. We deem them as natural because the perturbation process does not involve any artificial manipulation, in line with the definitions by Belinkov and Bisk (2018); Hendrycks et al. (2021); Pedraza et al. (2022); Agarwal et al. (2022); Le et al. (2022) (Figure 1). Results of robustness evaluation are therefore more representative of real-world applications. Similar to Belinkov and Bisk (2018), our approach utilises Wikipedia revision histories as the source of natural perturbations, given that the differences between revisions authentically capture the textual modifications made by human editors in the real world. Despite this, significant differences exist in the perturbation construction methodology between us. Perturbation in (Belinkov and Bisk, 2018) is restricted to single word replacements and applied on non-English source-side sentences in machine translation. In detail, they build a look-up table of possible lexical replacements by harvesting naturally occurring errors (typos, misspellings, etc.) from available corpora of French/German Wikipedia edits (Max and Wisniewski, 2010; Zesch, 2012). Afterwards, they replace every word in the source-side sentences with an error if one exists in the look-up table. Different from (Belinkov and Bisk, 2018), our approach does not restrict the perturbation level and utilise English Wikipedia. By comparing the variances between each adjacent revision, we identify perturbed versions for each Wikipedia reading passage in the original MRC benchmarks (if it exists). This enables us to capture more comprehensive and critical natural perturbation patterns (see Section 5.2) that can not be possible to capture in (Belinkov and Bisk, 2018). Our perturbation method only alter the reading context, while the questions and ground truth answers remain unchanged.
@@
With the established framework, we conduct extensive experiments on nine datasets, evaluating forty-two models, including recently proposed LLMs such as DeepSeek. Experimental results on Stanford Question Answering Dataset (SQuAD) (Rajpurkar et al., 2016, 2018) indicate that natural perturbations encompass rich linguistic variations and can lead to failures in the encoder-only models, while humans are almost undeterred by their presence. Crucially, these errors also transfer to larger and more powerful models, such as Flan-T5 and state-of-the-art (SOTA) LLMs. These findings also generalise to other and more challenging MRC benchmarks (e.g., DROP (Dua et al., 2019) and HotpotQA (Yang et al., 2018)) resulting in a decrease of SOTA LLMs’ performance, emphasising the harmful effects of natural perturbations. Adversarial re-training with either naturally or synthetically perturbed MRC instances can enhance the robustness of encoder-only models against natural perturbations, with the latter sometimes providing greater benefits. However, there is still ample room for improvement, calling for better defense strategies.
@@
A Wikipedia revision history-based framework to generate natural perturbed MRC benchmarks for realistic robustness evaluation.
Perturbed datasets for nine diverse MRC tasks. Two SQUAD challenge sets derived from error analysis of encoder-only models, on which SOTA LLMs struggle, even without being involved in the creation in any capacity.
Empirical demonstration of the validity of natural perturbations, their characterisation by different linguistic phenomena and their harmful effects on diverse model architectures across benchmarks generated with the proposed framework.
Showcasing adversarial re-training with natural or, especially, synthetic perturbations as a way to enhance the robustness of encoder-only MRC models against natural perturbations.
@@ 4.1 Datasets
We use nine English MRC datasets: SQuAD 1.1 (Rajpurkar et al., 2016), SQuAD 2.0 (Rajpurkar et al., 2018), BoolQ (Clark et al., 2019), DROP (Dua et al., 2019), HotpotQA (distractor) (Yang et al., 2018), TyDi QA (gold passage task in English) (Clark et al., 2020) and three datasets constructed in Adversarial QA (Bartolo et al., 2020): D(BiDAF), D(BERT) and D(RoBERTa). These are chosen as their reading passages are sourced from Wikipedia, thereby enabling the utilisation of Wikipedia editing histories to generate the naturally perturbed test set.
@@ 4.3 Evaluation Metrics
In line with existing literature, we choose the (instance-averaged) Token-F1 score to assess the performance of both encoder-only and encoder-decoder models (Rajpurkar et al., 2016), as on SQuAD-style test sets, they are optimised to output the shortest continuous span from the context as the answer (or predict the question as unanswerable) during inference. However, the outputs of the decoder-only models do not consistently adhere to the instruction due to their conversational style, rendering F1 unsuitable for evaluation. Consequently, we employ a more lenient metric, namely Inclusion Match (IM), which measures whether the response of the model contains any of the ground truth answers (Bhuiya et al., 2024). Furthermore, if the model’s output includes phrases such as “I cannot answer this/the question” or “unanswerable”333We collate a collection of such phrases by manually examining the decoder-only models’ outputs (Check Appendix D for the full set)., we deem that the model believes the question is not answerable. Model robustness is quantified by measuring the relative variation in performance (as reflected in the F1 or IM) under perturbations.
@@ 5 MRC under Natural Perturbation
In this section, we present and discuss the results of our experiments. We first evaluate encoder-only models on SQuAD to establish a baseline evaluation of model behaviour under natural perturbations. While neither represents the current SOTA, SQUAD’s simplicity and the stable, super-human performance of encoder-only models enable a focused and controlled examination of perturbation effects (Section 5.1), error sources (Section 5.2), and adversarial instance validity (Section 5.3). Then, we investigate the transferability of errors from encoder-only models to other architectures, showing both Flan-T5 and LLMs carry these errors significantly (Section 5.4) 444To supplement, we further evaluate the full test set on Flan-T5 and several LLMs, and measure the transferability of adversarial examples across all model architectures (Appendix E).. We finally generalise the findings from the baseline evaluation to SOTA LLMs and other more complex datasets (Section 5.5).
5.1 Are Encoder-only MRC Models Resilient to Natural Perturbation?
Table 1 presents the relative F1 change for all encoder-only MRC models on the naturally perturbed test set generated based on the SQuAD 1.1 and SQuAD 2.0 development set, respectively. It can be clearly seen from Table 1 that overall, the performance of all the examined models decreases, indicating that encoder-only MRC models suffer from natural perturbation. However, we notice that the performance drop of all models is negligible (the biggest drop is only %), which suggests that those models also exhibit considerable robustness to natural perturbations.
@@ 5.1 Are Encoder-only MRC Models Resilient to Natural Perturbation?
5.2 Error Analysis
Although encoder-only MRC models exhibit a relatively small performance gap, it remains worthwhile to investigate the sources of natural perturbation and reveal the perturbation phenomena contributing to models’ error. To this end, we manually label linguistic features between passages where models succeed and fail, to identify how they differ.
@@ 5.2 Error Analysis
5.3 Validity of Nature Adversarial Examples
@@ 5.4 Can Errors from Encoder-only Models Affect Other Architectures?
Table 2 shows the evaluation results on the newly generated challenge test sets. From the table, we observe that the errors caused by natural perturbation in encoder-only MRC models transfer to both Flan-T5 and LLMs. On the Nat_V1_Challenge, flan-t5-small demonstrates the greatest susceptibility to natural perturbations, experiencing a % decrease in F1; while among LLMs, Gemma-7B-IT emerges as the least robust, with a % IM drop, followed by Gemma-2B-IT (%) and Llama-3.1-8B-Instruct (%). Transitioning to the Nat_V2_Challenge, the base version of flan-t5 exhibits the largest performance decline at % and Falcon-7B-Instruct stands out as the LLM with the lowest robustness (%). Other LLMs such as Qwen2.5-7B-Instruct and deepseek-llm-7b-chat also show severe robustness loss, with drops of % and %, respectively. Further, we observe that the robustness of models under natural perturbations does not necessarily size-dependent. While larger models tend to exhibit greater robustness in some cases (e.g., Qwen2.5-14B-Instruct vs. Qwen2.5-3B-Instruct), exceptions within the Falcon and Llama model series suggest that factors beyond model size–such as pretraining data, training and fine-tuning methodology, and architectural differences may also significantly influence susceptibility to natural perturbations. In Appendix J, we showcase two adversarial examples targeting LLMs sourced from our generated challenge sets.
@@
5.5 Do Our Findings Generalise to Other MRC Datasets?
The two SQuAD datasets investigated previously are relatively simple, as they lack challenging features (Schlegel et al., 2020), leading to super-human performance of MRC models (Lan et al., 2020). To generalise our findings to more challenging MRC benchmarks, we apply the natural perturbation methodology (Section 3) to the development set of seven more datasets and assess the performance changes of multiple LLMs, as shown in Table 3. For DROP (Dua et al., 2019), we first use the GPT-4o mini to infer the likely Wikipedia article title from which each passage is retrieved 777Prompt: “Given a reading paragraph, return the Wikipedia page title from which it is likely retrieved.” and extract the revision histories for those articles. For HotpotQA (Yang et al., 2018), we only perturb the paragraphs containing the supporting facts.
| LLM | IM Relative Change (%) | ||||||
| BoolQ | DROP | HotpotQA | TYDI QA | D(BiDAF) | D(BERT) | D(RoBERTa) | |
| Gemma 2-2b-IT | |||||||
| Gemma 2-9b-IT | |||||||
| Llama-3.1-8B-Instruct | |||||||
| Llama-3.2-1B-Instruct | |||||||
| Llama-3.2-3B-Instruct | |||||||
| Mistral-7B-Instruct-v0.2 | |||||||
| OLMo-7B-0724-Instruct | |||||||
| Qwen2.5-3B-Instruct | |||||||
| Qwen2.5-7B-Instruct | |||||||
| Qwen2.5-14B-Instruct | |||||||
| Falcon3-7B-Instruct | |||||||
| Falcon3-10B-Instruct | |||||||
| deepseek-llm-7b-chat | |||||||
| average | |||||||
Overall, when natural perturbations are applied to more challenging benchmarks, SOTA LLMs also exhibit a lack of robustness. On average, the largest performance decline occurs on DROP (%), suggesting that natural perturbations significantly impair models’ discrete reasoning capabilities. Besides the DROP, the average performance degradation remains substantial across Adversarial QA and TyDi QA. HotpotQA, which requires multi-hop reasoning, also shows non-negligible degradation (%). These further demonstrate the broad and severe impact of natural perturbations on diverse MRC tasks.

### v2

@@ Abstract
As neural language models achieve human-comparable performance on Machine Reading Comprehension (MRC) and see widespread adoption, ensuring their robustness in real-world scenarios has become increasingly important. Current robustness evaluation research, though, primarily develops synthetic perturbation methods, leaving unclear how well they reflect real life scenarios. Considering this, we present a framework to automatically examine MRC models on naturally occurring textual perturbations, by replacing paragraph in MRC benchmarks with their counterparts based on available Wikipedia edit history. Such perturbation type is natural as its design does not stem from an arteficial generative process, inherently distinct from the previously investigated synthetic approaches. In a large-scale study encompassing SQuAD datasets and various model architectures we observe that natural perturbations result in performance degradation in pre-trained encoder language models. More worryingly, these state-of-the-art Flan-T5 and Large Language Models (LLMs) inherit these errors, with the largest observed drop reaching 28.28%. Further experiments demonstrate that our findings generalise to natural perturbations found in other more challenging MRC benchmarks such as DROP and HotpotQA. In an effort to mitigate these errors, we show that robustness to natural perturbations can be improved through adversarial training for encoder-only models or through in-context demonstrations of perturbed instances for LLMs, although a more generalisable and effective defence strategy remains to be developed.
@@ 1 Introduction
Transformer-based pre-trained language models demonstrate remarkable efficacy in addressing questions based on a given passage of text, a task commonly referred to as Machine Reading Comprehension (MRC) (Devlin et al., 2019; Brown et al., 2020; He et al., 2021; Wei et al., 2022; Touvron et al., 2023; OpenAI et al., 2024b). Despite these advancements, high-performing MRC systems are known to succeed by exploiting shortcuts in existing benchmark datasets rather than truly demonstrating understanding of the passage and question, thereby exhibiting a lack of robustness to various types of test-time perturbations. Such phenomenon is observed not only in encoder-only and encoder-decoder models (Ho et al., 2023; Schlegel et al., 2023), but also in state-of-the-art (SOTA) Large Language Models (LLMs) (Levy et al., 2023; Fang et al., 2023; Gupta et al., 2024).
Extensive textual perturbation methods have been developed to reveal the vulnerabilities of MRC models to various linguistic challenges (Ribeiro et al., 2018; Jiang and Bansal, 2019; Welbl et al., 2020; Tan et al., 2020; Tan and Joty, 2021; Schlegel et al., 2021; Cao et al., 2022; Tran et al., 2023). Despite the insights they provide, all of them craft perturbations in a synthetic manner—that is, based on hypothesised manipulation strategies (Figure 1 Top)—which may not accurately reflect the types of challenges encountered in real-world scenarios. As a result, there is a risk of overlooking the genuine weaknesses of reading comprehension systems when deployed in practical settings, potentially hindering efforts to improve their reliability in real applications.
@@
To counteract this issue, in this paper, we develop a framework to inject textual changes that arise in real-world conditions into MRC datasets and audit how well contemporary language models perform under such perturbations. We deem them as natural because the perturbation process does not involve any artificial manipulation, in line with the definitions by Belinkov and Bisk (2018); Hendrycks et al. (2021); Pedraza et al. (2022); Agarwal et al. (2022); Le et al. (2022) (Figure 1 Bottom). Results of robustness evaluation are therefore more representative of real-world applications. Similar to Belinkov and Bisk (2018), our approach utilises Wikipedia revision histories as the source of natural perturbations, given that the differences between revisions authentically capture the textual modifications made by human editors in the real world. Despite this, significant differences exist in the perturbation construction methodology between us. Perturbation in (Belinkov and Bisk, 2018) is restricted to single word replacements and applied on non-English source-side sentences in machine translation. In detail, they build a look-up table of possible lexical replacements by harvesting naturally occurring errors (typos, misspellings, etc.) from available corpora of French/German Wikipedia edits (Max and Wisniewski, 2010; Zesch, 2012). Afterwards, they replace every word in the source-side sentences with an error if one exists in the look-up table. Different from (Belinkov and Bisk, 2018), our approach does not restrict the perturbation level and utilise English Wikipedia. By comparing the variances between each adjacent revision, we identify perturbed versions for each Wikipedia reading passage in the original MRC benchmarks (if it exists). This enables us to capture more comprehensive and critical natural perturbation patterns (see Section 5.2) that can not be possible to capture in (Belinkov and Bisk, 2018). Our perturbation method only alter the reading context, while the questions and ground truth answers remain unchanged.
@@
With the established framework, we conduct extensive experiments on six datasets, evaluating forty-two models, including recently proposed LLMs. Experimental results on Stanford Question Answering Dataset (SQuAD) (Rajpurkar et al., 2016, 2018) indicate that natural perturbations encompass rich linguistic variations and can lead to failures in the encoder-only models, while humans are almost undeterred by their presence. Crucially, these errors also transfer to larger and more powerful models, such as Flan-T5 and SOTA LLMs, with performance drops ranging from % to %. These findings also generalise to other and more challenging MRC benchmarks (e.g., Mistral-7B-Instruct-v0.2’s % decrease on BoolQ (Clark et al., 2019) and Llama-3.2-1B-Instruct’s % decline on DROP (Dua et al., 2019)), emphasising the harmful effects of natural perturbations. Adversarial re-training/in-context demonstration with either naturally or synthetically perturbed MRC instances can enhance the robustness against natural perturbations, with the latter sometimes providing greater benefits. However, there is still ample room for improvement, calling for better defense strategies.
@@
A framework—based on Wikipedia revision history—for studying model robustness under real-world natural perturbations. This is relevant, even in the LLM era, as our framework can be applied to any other tasks with input from Wikipedia and also to any types of models.
Perturbed datasets for six diverse MRC tasks. Two SQUAD challenge sets derived from error analysis of encoder-only models, on which SOTA LLMs struggle, even without being involved in the creation in any capacity.
Empirical demonstration of the validity of natural perturbations across both encoder-only models and LLMs, their characterisation by different linguistic phenomena and their harmful effects on diverse model architectures across benchmarks generated with the proposed framework.
Showcasing adversarial re-training with natural or, especially, synthetic perturbations, as well as adversarial in-context demonstrations as a way to enhance the robustness of encoder-only models and LLMs, respectively, against natural perturbations.
@@ 4.1 Datasets
We use six English MRC datasets: SQuAD 1.1 (Rajpurkar et al., 2016), SQuAD 2.0 (Rajpurkar et al., 2018), BoolQ (Clark et al., 2019), DROP (Dua et al., 2019), HotpotQA (distractor) (Yang et al., 2018) and TyDi QA (gold passage task in English) (Clark et al., 2020). These are chosen as their reading passages are sourced from Wikipedia, thereby enabling the utilisation of Wikipedia editing histories to generate the naturally perturbed test set.
@@ 4.3 Evaluation Metrics
In line with existing literature, we choose the (instance-averaged) Token-F1 score to assess the performance of both encoder-only and encoder-decoder models (Rajpurkar et al., 2016), as on SQuAD-style test sets, they are optimised to output the shortest continuous span from the context as the answer (or predict the question as unanswerable) during inference. However, the outputs of the decoder-only models do not consistently adhere to the instruction due to their conversational style, rendering F1 unsuitable for evaluation. Consequently, we employ a more lenient metric, namely Inclusion Match (IM), which measures whether the response of the model contains any of the ground truth answers (Bhuiya et al., 2024). Furthermore, if the model’s output includes phrases such as “I cannot answer this/the question” or “unanswerable”333We collate a collection of such phrases by manually examining the decoder-only models’ outputs (Check Appendix D for the full set)., we deem that the model believes the question is not answerable. Model robustness is quantified by measuring the relative variation in performance (as reflected in the F1 or IM) under natural perturbations.
@@ 5 MRC under Natural Perturbation
In this section, we present and discuss the results of our experiments. We first evaluate encoder-only models on SQuAD to establish a baseline evaluation of model behaviour under natural perturbations. While neither represents the current SOTA, SQUAD’s simplicity, the stable, super-human performance of encoder-only models, and it’s advantage of being free from benchmark leakage impact enable a focused and controlled examination of perturbation effects (Section 5.1), error sources (Section 5.2), and adversarial instance validity (Section 5.3). Then, we investigate the transferability of errors from encoder-only models to other architectures, showing both Flan-T5 and LLMs carry these errors significantly (Section 5.4) 444To supplement, we further evaluate the full test set on Flan-T5 and several LLMs, and measure the transferability of adversarial examples across all model architectures (Appendix E).. We finally generalise the findings from the baseline evaluation to SOTA LLMs and other more complex datasets (Section 5.5).
5.1 Are Encoder-only Models Resilient to Natural Perturbation?
Table 1 presents the relative F1 change for encoder-only MRC models on the naturally perturbed SQuAD test set. It can be seen that overall, the performance of all the examined models decreases, indicating that encoder-only MRC models suffer from natural perturbation. Nonetheless, the performance decline of all models is not at a significant level, with the biggest drop being only %. This suggests that these models also exhibit considerable robustness to natural perturbation.
@@ 5.1 Are Encoder-only Models Resilient to Natural Perturbation?
5.2 Which Categories of Natural Perturbation Lead to Model Failure?
To investigate the sources of natural perturbation and reveal the perturbation phenomena contributing to encoder-only models’ error, we manually label linguistic features between passages where models succeed and fail, to identify how they differ.
@@ 5.2 Which Categories of Natural Perturbation Lead to Model Failure?
5.3 To What Extent Do Natural Adversarial Examples Preserve Validity?
@@ 5.4 Can Errors from Encoder-only Models Affect Other Architectures?
Table 2 shows the evaluation results on the newly generated challenge test sets. From the table, we observe that the errors caused by natural perturbation in encoder-only MRC models transfer to both Flan-T5 and LLMs. On the Nat_V1_Challenge, flan-t5-small demonstrates the greatest susceptibility to natural perturbations, experiencing a % decrease in F1; while among LLMs, Gemma-7B-IT emerges as the least robust, with a % IM drop, followed by Gemma-2B-IT (%) and Llama-3.1-8B-Instruct (%). Transitioning to the Nat_V2_Challenge, the base version of flan-t5 exhibits the largest performance decline at % and Falcon-7B-Instruct stands out as the LLM with the lowest robustness (%). Other LLMs such as Qwen2.5-7B-Instruct and deepseek-llm-7b-chat also show severe robustness loss, with drops of % and %, respectively. Further, we observe that the robustness of models under natural perturbations does not necessarily size-dependent. While larger models tend to exhibit greater robustness in some cases (e.g., Qwen2.5-14B-Instruct vs. Qwen2.5-3B-Instruct), exceptions within the Falcon and Llama model series suggest that factors beyond model size–such as training corpora, training and fine-tuning methodology, and architectural differences may also significantly affect susceptibility to natural perturbations. In Appendix J, we showcase two adversarial examples targeting LLMs sourced from our generated challenge sets.
@@
5.5 Do Our Findings Generalise to Other Challenging MRC Datasets?
The two SQuAD datasets investigated previously are relatively simple, as they lack challenging features (Schlegel et al., 2020), leading to super-human performance of MRC models (Lan et al., 2020). To generalise our findings to more challenging MRC benchmarks, we apply the natural perturbation methodology (Section 3) to the development set of four more datasets and assess the performance changes of multiple LLMs, as shown in Table 3. For DROP (Dua et al., 2019), we first use the GPT-4o mini to infer the likely Wikipedia article title from which each passage is retrieved 777This is because the raw Wikipedia title information cannot be found in the original development set of DROP. We use the prompt: “Given a reading paragraph, return the Wikipedia page title from which it is likely retrieved.” and extract the revision histories for those articles. For HotpotQA (Yang et al., 2018), we only perturb the paragraphs containing the supporting facts, while the distracting passages remain unchanged. Furthermore, using the validity verification method described in Section 5.3, we manually verify the validity percentage of all adversarial examples in DROP and TyDi QA, as well as randomly selected adversarial examples from BoolQ and HotpotQA, as reported in Table 3.
| LLM | IM Relative Change (%) | |||
| BoolQ | DROP | HotpotQA | TYDI QA | |
| adversarial validity (%) (Cohen’s ) | (0.54) | (0.46) | (0.6) | (0.52) |
| Gemma 2-2b-IT | ||||
| Gemma 2-9b-IT | ||||
| Llama-3.1-8B-Instruct | ||||
| Llama-3.2-1B-Instruct | ||||
| Llama-3.2-3B-Instruct | ||||
| Mistral-7B-Instruct-v0.2 | ||||
| OLMo-7B-0724-Instruct | ||||
| Qwen2.5-7B-Instruct | ||||
| Qwen2.5-14B-Instruct | ||||
| Falcon3-7B-Instruct | ||||
| Falcon3-10B-Instruct | ||||
Overall, when natural perturbations are applied to other more challenging benchmarks, SOTA LLMs also exhibit a lack of robustness, with the largest performance decrease observed for Llama-3.2-1B-Instruct on DROP. This further demonstrates the broad and severe impact of natural perturbations on diverse MRC tasks, especially in light of on three out of four benchmarks, our human annotators are still able to correctly answer over % of the adversarial examples. BoolQ exhibits a lower adversarial validity rate (%). However, in most cases, this is not due to the perturbation degrading the passage, but rather the poor quality of the original instance (Tedeschi et al., 2023), i.e., even with the corresponding original passage, humans are unable to assign the correct label. For instance, from the annotators’ perspective, the question itself may be ambiguous, such as “do you need a visa to visit oman”, or entirely unanswerable due to missing information in the original passage. Our human annotations also observe diverse error patterns in LLMs caused by natural perturbations (e.g., Copy Editing, Elaboration and Vandalism), as what are presented in Figure 3. Finally, we note that for a very small proportion of adversarial examples, the limitations of the IM metric mean that the model’s answer under perturbation does not necessarily indicate an error, emphasising the need for future efforts to address this concern.